OpenStudy (anonymous):

Chemistry Chromatography help please? Will fan+medal! Help finding how far the Ds and Df moved?

OpenStudy (anonymous):

Ds means the distance traveled by compounds and Dr is distance traveled by solvent I cannot for the life of me read the ruler or figure this out

OpenStudy (anonymous):

I think the Df is 0 and the Ds is 0.4cm but that wouldn't make sense because then Rf wouldn't equal anything and the Rf is the Ds/Df

OpenStudy (anonymous):

*Df not Dr sorry

OpenStudy (cuanchi):

In the picture where it said Ds looks like that is were the sample was loaded in the paper (this is not Ds) .And where it said Df is were the solvent got after the run. It is not very clear that you have a spot at 4cm but if you have one that will be the Ds. then the Rf= Ds/Df Ds the distance between the load sample to the spot = 4cm Df the distance between the load sample to the end of the solvent run = 11.6 cm Rf= 4/11.6 = 0.35 your sample migrate ~ 1/3 of the total run.

OpenStudy (anonymous):

Oh my gosh thank you so much!!! I'm so terrible at the whole ruler measuring. Thank you!

OpenStudy (anonymous):

May I ask how you found the answers? There are a few more questions just like that one but I don't want to bother anyone.

OpenStudy (anonymous):

@Cuanchi ?

OpenStudy (anonymous):

Nevermind I will just post them in a new question